# Flow-Attentional Graph Neural Networks

### v1

@@ Abstract
Keywords Graph Neural Networks  $\cdot$ Directed Acyclic Graphs  $\cdot$ Electronic Circuits  $\cdot$ Power Grid
@@ 1 Introduction
Graph neural networks (GNNs) (Scarselli et al., 2008; Kipf and Welling, 2017) have emerged as a powerful framework that extends the scope of deep learning from Euclidean to graph-structured data, which is prevalent across many real-world domains, such as social networks (Fan et al., 2019), recommender systems (Wu et al., 2022), materials science (Reiser et al., 2022) or epidemiology (Liu et al., 2024). Especially attention-based GNNs have become increasingly popular due to their ability to select relevant features adaptively (Sun et al., 2023). As graph data becomes increasingly common, advancing GNN architectures is crucial for improving performance in tasks such as node classification (Hamilton et al., 2017), graph regression (Gilmer et al., 2017), or link prediction (Zhang and Chen, 2018).
In many important applications of GNNs, graphs are naturally associated with a flow of physical resources, such as electrical current in electronic circuits (Sánchez et al., 2023) or power grids (Liao et al., 2021), traffic in transportation networks (Jiang and Luo, 2022), water in river networks (Sun et al., 2021), or raw materials and goods in supply chains (Kosasih and Brintrup, 2022). All nodes in these resource flow graphs, except for source and target nodes, are subject to Kirchhoff’s first law, which states that the sum of all incoming and outgoing flows must be zero, reflecting the conservation of resources. By contrast, informational graphs—such as computation graphs, social networks, or citation networks—are not associated with any physical flow but rather represent relationships or information transfer. Information can be arbitrarily duplicated in these graphs, unlike in flow graphs, where such duplication would violate the conservation law.
As a result, two non-isomorphic graphs may be equivalent as informational graphs but non-equivalent as flow graphs. For example, in a computational graph, the output of a sine operation can be freely duplicated. Thus, the two non-isomorphic graph structures in Fig. 1a represent the same computation. However, the same graph structures may also represent electronic circuits governed by Kirchhoff’s first law (see Fig. 1b). In this case, the two circuits are different because combining or splitting resistors would change their electrical properties.
@@
Main Contributions. Inspired by the conservation law in resource flow graphs, we propose flow attention on graphs, which normalizes attention scores across outgoing neighbors instead of incoming ones. This simple but effective modification can be applied to any attention-based GNN and allows the model to better capture the physical flow of a graph. We discuss the expressivity of the resulting models and demonstrate that flow attention enables the discrimination of any DAG from its computation tree. Based on this observation, we propose FlowDAGNN, a flow-attentional GNN for DAGs. Finally, we conduct extensive experiments on multiple datasets, including cascading failure analysis on power grids and property prediction on electronic circuits, covering undirected graphs and DAGs. Our results indicate that flow attention improves the performance of attention-based GNNs on graph-level classification and regression tasks.  [^1] The code is available at https://github.com/pasplett/FlowGNN
@@ 2 Related Work
In recent years, many new GNN models have been specifically designed for different graph types (Thomas et al., 2023). Despite their fundamental differences, informational graphs and flow graphs are mainly treated by the same basic message-passing layers, such as GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018) or GIN (Xu et al., 2019). In these models, messages exchanged between neighboring nodes do not depend on the number of recipients. Instead, the information is arbitrarily duplicated and passed to all neighbors. GCN (Kipf and Welling, 2017) applies a symmetric neighborhood normalization but exhibits limited expressivity and performance. Attention-based GNNs, such as GAT, GATv2 (Brody et al., 2022) or Graph Transformer (Shi et al., 2021) adaptively weight neighboring node features leading to improved representation learning. However, the attention weights are obtained through normalization across incoming neighbors allowing for arbitrary message duplication. Our approach normalizes across outgoing neighbors instead, which better captures the physical flow of a graph while preserving the benefits of graph attention.
Many flow graphs, including the example graphs in Fig. 1, can be naturally expressed as DAGs, e.g., operational amplifiers (Dong et al., 2023) or material flow networks (Perera et al., 2018). Before GNNs were introduced, recursive neural networks were applied to DAGs (Sperduti and Starita, 1997; Frasconi et al., 1998) and contextual recursive cascade correlation was proposed to overcome limitations in expressivity (Hammer et al., 2005). However, these early works lacked the advantages of modern GNNs, which have been extended to DAGs in recent years (Zhang et al., 2019; Thost and Chen, 2021). In directed acyclic GNNs, nodes are typically updated sequentially following the partial order of the DAG, and the final target node representations are used as the graph embedding. While these sequential models outperform undirected GNNs on DAG datasets, they still aggregate node neighborhood information similarly. Therefore, they only encode the computation tree of the output nodes and not the exact structure of the DAG, resulting in limited expressivity.
@@ 3 Preliminaries
Computation Tree. Let $\mathcal{D}=(\mathcal{V},\mathcal{E},r)$ be a rooted DAG with a unique final node called the root $r$. Its computation tree $\gamma(\mathcal{D})$ is obtained by leaving each node $v$ with at most one successor, and for any node $v$ with $n\geq 2$ successors, replacing $v$ by $n$ copies $v_{1},\dots,v_{n}$, each connected to exactly one of $v$’s successors and inheriting all of $v$’s incoming edges. This procedure yields a rooted tree with the same root $r$. See App. D for a visualization.
@@
$\bm{h}_{i}^{\prime}=\phi\left(\bm{h}_{i},\underset{j\in\mathcal{N}_{\text{in}}(i)}{\bigoplus}f\left(\bm{h}_{j}\right)\right),$ (2)
with a learnable message function $f$, an aggregator $\oplus$, e.g., sum or mean, and an update function $\phi$. The choice of $\phi$, $\oplus$, and $f$ are defining the design of a specific GNN model.
@@
$\bm{h}_{i}^{\prime}=\phi\left(\bm{h}_{i},\underset{j\in\mathcal{N}_{\text{in}}(i)}{\bigoplus}f\left(\bm{h}_{j}^{\prime}\right)\right).$ (3)
The most widely used directed acyclic GNNs are D-VAE (Zhang et al., 2019) and DAGNN (Thost and Chen, 2021), which uses standard attention for aggregation. Both models utilize gated recurrent units (GRU) as the update function $\phi$ and are briefly explained in App. B. As an alternative to sequential models, DAGs can be encoded using Transformer-based architectures, such as PACE (Dong et al., 2022) or DAGformer (Luo et al., 2023).
4 Attention is not all you need for Flow Graphs
@@ 4.1 Attentional GNNs
Popular attentional GNNs include GAT (Veličković et al., 2018), GATv2 (Brody et al., 2022) and Graph Transformer (GT) (Shi et al., 2021), which mainly differ in the choice of the scoring function $e$ (see App. C). The graph attention mechanism is visualized in Fig. 2a.
@@ 4.2 Attention on DAGs
A directed acyclic GNN sequentially updates each node’s feature vector until it arrives at one or more final nodes. If there are multiple final nodes, their representations can be gathered into a single virtual node. Since all other nodes are ancestors of this final node, its representation can be used to characterize the whole DAG. The computation history of the final node representation can be visualized as a rooted subtree, which means that if two non-isomorphic DAGs represent the same computation, the rooted subtree structures of their final node representations are equivalent. Therefore, since standard attention weights only depend on incoming neighborhoods, directed acyclic GNNs with standard attention mechanisms cannot discriminate a DAG $\mathcal{D}$ from its computation tree $\gamma(\mathcal{D})$ (see Fig. 3 in App. D).
@@ 5.1 Flow-Attentional GNNs
Standard attention mechanisms normalize attention scores across all incoming edges. Therefore, a message does not depend on how many nodes it is forwarded to, which means it can be duplicated arbitrarily, contradicting the resource conservation concept inherent in flow graphs. Therefore, we propose an alternative graph attention mechanism that normalizes the attention scores across outgoing edges instead (see Fig. 2b). We denote the resulting flow attention weights as $\beta_{ij}$ to distinguish them from the standard attention weights $\alpha_{ij}$:
@@
However, since the messages are multiplied with the flow attention weights $\beta_{ij}$, they now also depend on the neighborhood of the message’s sender, i.e., node $j$. In this way, we ensure that a message transmitted by any node cannot be duplicated arbitrarily but instead is distributed among all outgoing neighbors. We define a flow-attentional graph neural network (FlowGNN) as a modified version of an attentional GNN, which utilizes the flow attention mechanism from Eq. 9 for aggregating node neighborhood information instead of standard attention. Furthermore, we denote the corresponding FlowGNN versions of standard attentional GNNs as FlowGAT, FlowGATv2, FlowGT, etc.
@@
Thereby, $\psi_{0}(j,i)$ is the absolute outgoing flow from a source or target node. Then $\psi_{\beta}$ is a flow on $\mathcal{G}$ that satisfies Kichhoff’s first law.
@@
Assume $\mathcal{X}_{1}=(S,m)$ and $\mathcal{X}_{2}=(S,k\cdot m)$ are multisets with the same distribution of elements for some $k\in\mathbb{N}_{\geq 2}$. Furthermore, assume that all nodes $s\in S$ have the same outgoing neighborhood $\mathcal{N}_{\text{out}}(s)$. Then $\exists~{}\phi,f$ such that $\bm{h}_{\text{flow}}^{\prime}(\mathcal{X}_{1})\neq\displaystyle\bm{h}_{\text{flow}}^{\prime}(\mathcal{X}_{2})$.
Therefore, flow-attentional GNNs are particularly well-suited for tasks where not only statistical information, but also the precise graph structure, plays a crucial role. This is especially true when features occur repeatedly, e.g., in the case of multiple identical resistors in an electronic circuit.
@@ 5.2 Flow-Attention on DAGs
For every DAG $\mathcal{D}\in\Delta$ with $\mathcal{D}\notin\mathcal{T}$ (i.e., a true DAG) and its computation tree $\gamma(\mathcal{D})$ there exists a flow-attentional directed acyclic GNN $\mathcal{A}_{\text{flow}}$ such that:
@@
for the graph $\mathcal{D}$ and its computation tree $\gamma(\mathcal{D})$, respectively. Since $\mathcal{D}$ is not a tree, there is at least one node $j$ with more than one successor; hence there is a predecessor $j$ for which $\beta_{ij}<1$. We then need to show that $\bm{h}^{\prime}_{i}\neq\tilde{\bm{h}}^{\prime}_{i}$:
@@
Hence, with choosing $\phi$ to be injective, $\mathcal{A}_{\text{flow}}$ distinguishes $\mathcal{D}$ from $\gamma(\mathcal{D})$. ∎
From Theorem 5.4 we conclude that a flow-attentional directed acyclic GNN can discriminate the example DAGs from Fig. 1 for the right choice of $\phi$ and $f$. In practice, we can model the composition $f^{(l)}\circ\phi^{(l-1)}$ on the $l$-th GNN layer by a universal approximator, e.g., GRU (Schäfer and Zimmermann, 2006; hoon Song et al., 2023).
@@ 5.3 FlowDAGNN
We propose a flow-attentional version of the attention-based DAGNN (Thost and Chen, 2021). A naive approach would be to simply replace the attention weights $\alpha_{ij}$ with flow attention weights $\beta_{ij}$. Due to the sequential nature of directed acyclic GNNs, the computation of the flow attention weight $\beta_{ij}$ in a DAG only depends on the node $i$ and all its ancestors. However, in contrast to standard attention weights, flow attention weights are forward-directed, which means that they should also be conditioned on all descendants of the node $i$. Analogously, the electrical current splitting up from one node into multiple branches depends on the whole branch and not only on the first node in each branch. In App. E, we give a simple example of an electronic circuit illustrating this situation.
Hence, we construct a FlowDAGNN layer from two sublayers (see Fig. 2c). In the first sublayer (we call it the reverse pass), we invert all edges of the DAG $\mathcal{D}$ and apply a standard DAGNN layer to the reverse DAG $\tilde{\mathcal{D}}$. This is equivalent to performing the aggregation over all successor nodes in the original DAG $\mathcal{D}$ instead of over all predecessors:
$\bm{m}_{i}^{\text{rv}}$ $=\sum_{j\in\mathcal{N}_{out}(i)}\alpha_{ij}\left(\bm{h}_{i},\bm{h}_{j}^{\text{rv}}\right)\bm{h}_{j}^{\text{rv}},$ (12)
$\alpha_{ij}\left(\bm{h}_{i},\bm{h}_{j}^{\text{rv}}\right)$ $=\underset{j\in\mathcal{N}_{out}(i)}{\text{softmax}}\left((\bm{w}_{1}^{\text{rv}})^{\text{T}}\bm{h}_{i}+(\bm{w}_{2}^{\text{rv}})^{\text{T}}\bm{h}_{j}^{\text{rv}}\right),$ (13)
$\bm{h}_{i}^{\prime}$ $=\bm{h}_{i}^{\text{rv}}=\text{GRU}(\bm{h}_{i},\bm{m}_{i}^{\text{rv}}).$ (14)
In the second sublayer, we perform a forward pass on the original DAG $\mathcal{G}$. However, this time we are applying the flow attention mechanism described in Section 5.1 to compute flow attention weights $\beta_{ij}$:
$\bm{m}_{i}^{\text{fw}}$ $=\sum_{j\in\mathcal{N}_{out}(i)}\beta_{ij}\left(\bm{h}_{i}^{\text{rv}},\bm{h}_{j}^{\text{fw}}\right)\bm{h}_{j}^{\text{fw}},$ (15)
$\beta_{ij}\left(\bm{h}_{i}^{\text{rv}},\bm{h}_{j}^{\text{fw}}\right)$ $=\underset{i\in\mathcal{N}_{out}(j)}{\text{softmax}}\left((\bm{w}_{1}^{\text{fw}})^{\text{T}}\bm{h}_{i}^{\text{rv}}+(\bm{w}_{2}^{\text{fw}})^{\text{T}}\bm{h}_{j}^{\text{fw}}\right),$ (16)
$\bm{h}_{i}^{\text{fw}}$ $=\text{GRU}(\bm{h}_{i}^{\text{rv}},\bm{m}_{i}^{\text{fw}}).$ (17)
Since the hidden states $\bm{h}_{i}^{\text{rv}}$ of the reverse pass contain information about all descendants of the node $i$, and the hidden states $\bm{h}_{j}^{\text{fw}}$ contain information about all ancestors of the node $j$, the computation of the flow attention weights $\beta_{ij}$ essentially takes into account information about all nodes of the graph that are connected to the node $i$.
After $L$ FlowDAGNN layers, we compute the graph-level representation from both the reverse pass representations of the start nodes as well as the forward pass representations of the end nodes and concatenate across layers:
[…]

### v2

@@ 1 Introduction
Graph neural networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2017) have emerged as a powerful framework that extends the scope of deep learning from Euclidean to graph-structured data, which is prevalent across many real-world domains, such as social networks (Fan et al., 2019), recommender systems (Wu et al., 2022), materials science (Reiser et al., 2022), or epidemiology (Liu et al., 2024). Especially attention-based GNNs have become increasingly popular due to their ability to select relevant features adaptively (Sun et al., 2023). As graph data becomes increasingly common, advancing GNN architectures is crucial for improving performance in tasks such as node classification (Hamilton et al., 2017), graph regression (Gilmer et al., 2017), or link prediction (Zhang & Chen, 2018).
In many important applications of GNNs, graphs are naturally associated with a flow of physical resources, such as electrical current in electronic circuits (Sánchez et al., 2023) or power grids (Liao et al., 2021), traffic in transportation networks (Jiang & Luo, 2022), water in river networks (Sun et al., 2021), or raw materials and goods in supply chains (Kosasih & Brintrup, 2022). All nodes in these resource flow graphs, except for source and target nodes, are subject to Kirchhoff’s first law, which states that the sum of all incoming and outgoing flows must be zero, reflecting the conservation of resources. By contrast, informational graphs—such as computation graphs, social networks, or citation networks—are not associated with any physical flow but rather represent relationships or information transfer. Information can be arbitrarily duplicated in these graphs, unlike in flow graphs, where such duplication would violate the conservation law.
As a result, two non-isomorphic graphs may be equivalent as informational graphs but non-equivalent as flow graphs. For example, in a computational graph, the output of a sine operation can be freely duplicated. Thus, the two non-isomorphic graph structures in Fig. 1a represent the same computation. However, the same graph structures may also represent electronic circuits governed by Kirchhoff’s first law (see Fig. 1b). In this case, the two circuits are different because combining or splitting resistors changes their electrical properties.
@@
Main Contributions. Inspired by the conservation law in resource flow graphs, we propose flow attention on graphs, which normalizes attention scores across outgoing neighbors instead of incoming ones. This simple but effective modification can be applied to any attention-based GNN and allows the model to better capture the physical flow of a graph. We discuss the expressivity of the resulting models and demonstrate that flow attention enables the discrimination of any DAG from its computation tree. Based on this observation, we propose FlowDAGNN, a flow-attentional GNN for DAGs. Finally, we conduct extensive experiments on multiple datasets, including cascading failure analysis on power grids and property prediction on electronic circuits, covering undirected graphs and DAGs. Our results indicate that flow attention improves the performance of attention-based GNNs on graph-level classification and regression tasks.  [^1] The code is available at https://github.com/pasplett/FlowGNN.
@@ 2 Related Work
In recent years, many new GNN models have been specifically designed for different graph types (Thomas et al., 2023). Despite their fundamental differences, informational graphs and flow graphs are mainly treated by the same basic message-passing layers, such as GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), or GIN (Xu et al., 2019). In these models, messages exchanged between neighboring nodes do not depend on the number of recipients. Instead, the information is arbitrarily duplicated and passed to all neighbors. GCN (Kipf & Welling, 2017) applies a symmetric neighborhood normalization but exhibits limited expressivity and performance. Attention-based GNNs, such as GAT, GATv2 (Brody et al., 2022), or UniMP (Shi et al., 2021), adaptively weight neighboring node features, leading to improved representation learning. However, the attention weights are obtained through normalization across incoming neighbors, allowing for arbitrary message duplication. Our approach normalizes across outgoing neighbors instead, which better captures the physical flow of a graph while preserving the benefits of graph attention.
Many flow graphs, including the example graphs in Fig. 1, can be naturally expressed as DAGs, e.g., operational amplifiers (Dong et al., 2023) or material flow networks (Perera et al., 2018). Before GNNs were introduced, recursive neural networks were applied to DAGs (Sperduti & Starita, 1997; Frasconi et al., 1998), and contextual recursive cascade correlation was proposed to overcome limitations in expressivity (Hammer et al., 2005). However, these early works lacked the advantages of modern GNNs, which have been extended to DAGs in recent years (Zhang et al., 2019; Thost & Chen, 2021). In directed acyclic GNNs, nodes are typically updated sequentially following the partial order of the DAG, and the final target node representations are used as the graph embedding. Although these sequential models outperform undirected GNNs on DAG datasets, they still aggregate node neighborhood information similarly. This means they only encode the computation tree of the output nodes and not the exact structure of the DAG, resulting in limited expressivity.
@@ 3 Preliminaries
Computation Tree. Let $\mathcal{D}=(\mathcal{V},\mathcal{E},r)$ be a rooted DAG with a unique final node called the root $r$. Its computation tree $\gamma(\mathcal{D})$ is obtained by leaving each node $v$ with at most one successor, and for any node $v$ with $n\geq 2$ successors, replacing $v$ by $n$ copies $v_{1},\dots,v_{n}$, each connected to exactly one of $v$’s successors and inheriting all of $v$’s incoming edges. This procedure yields a rooted tree with the same root $r$. See App. LABEL:app:dag_tree for a visualization.
@@
j∈Nin(i)⨁f(hj)), (2)
with a learnable message function $f$, an aggregator $\oplus$, e.g., sum or mean, and an update function $\phi$. The choice of $\phi$, $\oplus$, and $f$ defines the design of a specific GNN model.
@@
j∈Nin(i)⨁f(hj′)). (3)
The most widely used directed acyclic GNNs are D-VAE (Zhang et al., 2019) and DAGNN (Thost & Chen, 2021), the latter of which uses standard attention for aggregation. Both models utilize gated recurrent units (GRU) as the update function $\phi$ and are briefly explained in App. B. As an alternative to sequential models, DAGs can be encoded using Transformer-based architectures, such as PACE (Dong et al., 2022) or DAGformer (Luo et al., 2023).
4 Graph Attention and its Limitations
@@ 4.1 Attentional GNNs
Popular attentional GNNs include GAT (Veličković et al., 2018), GATv2 (Brody et al., 2022), and TransformerConv (TC)  [^2] Please note that TransformerConv is not a Graph Transformer with a global attention module (such as GraphGPS (Rampášek et al., 2022)) but a message-passing GNN with a local attention mechanism adopted from the vanilla Transformer (Vaswani et al., 2017). It was originally introduced under the name Graph Transformer as part of the UniMP model (Shi et al., 2021). (Shi et al., 2021), which mainly differ in the choice of the scoring function $e$ (see App. LABEL:app:scoring). The graph attention mechanism is visualized in Fig. 2a.
@@ 4.2 Attention on DAGs
A directed acyclic GNN sequentially updates each node’s feature vector until it arrives at one or more final nodes. If there are multiple final nodes, their representations can be gathered into a single virtual node. Since all other nodes are ancestors of this final node, its representation can be used to characterize the whole DAG. The computation history of the final node representation can be visualized as a rooted subtree, which means that if two non-isomorphic DAGs represent the same computation, the rooted subtree structures of their final node representations are equivalent. Therefore, since standard attention weights only depend on incoming neighborhoods, directed acyclic GNNs with standard attention mechanisms cannot discriminate a DAG $\mathcal{D}$ from its computation tree $\gamma(\mathcal{D})$ (see Fig. LABEL:fig:dag_tree in App. LABEL:app:dag_tree).
@@ 5.1 Flow-Attentional GNNs
Standard graph attention mechanisms normalize attention scores across all incoming edges. Therefore, a message does not depend on the number of nodes it is forwarded to and can be duplicated arbitrarily, contradicting the resource conservation concept inherent in flow graphs. Therefore, we propose an alternative graph attention mechanism that normalizes the attention scores across outgoing edges instead (see Fig. 2b). We denote the resulting flow attention weights as $\beta_{ij}$ to distinguish them from the standard attention weights $\alpha_{ij}$:
@@
However, since the messages are multiplied with the flow attention weights $\beta_{ij}$, they now also depend on the neighborhood of the message’s sender, i.e., node $j$. In this way, we ensure that a message transmitted by any node cannot be duplicated arbitrarily but instead is distributed among all outgoing neighbors. We define a flow-attentional graph neural network (FlowGNN) as a modified version of an attentional GNN, which utilizes the flow attention mechanism from Eq. 9 for aggregating node neighborhood information instead of standard attention. Furthermore, we denote the corresponding FlowGNN versions of standard attentional GNNs as FlowGAT, FlowGATv2, FlowTC, etc.
@@
Thereby, $\psi_{0}(j,i)$ is the absolute outgoing flow from a source or target node. Then $\psi_{\beta}$ is a flow on $\mathcal{G}$ that satisfies Kirchhoff’s first law.
@@
Assume $\mathcal{X}_{1}=(S,m)$ and $\mathcal{X}_{2}=(S,k\cdot m)$ are multisets with the same distribution of elements for some $k\in\mathbb{N}_{\geq 2}$. Furthermore, assume that all nodes $s\in S$ have the same outgoing neighborhood $\mathcal{N}_{\text{out}}(s)$ in $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$. Then $\exists~\phi,f$ such that $\bm{h}_{\text{flow}}^{\prime}(\mathcal{X}_{1})\neq\displaystyle\bm{h}_{\text{flow}}^{\prime}(\mathcal{X}_{2})$.
Therefore, flow-attentional GNNs are particularly well-suited for tasks where not only statistical information but also the precise graph structure plays a crucial role. This is especially true when features occur repeatedly, e.g., in the case of multiple identical resistors in an electronic circuit.
@@ 5.2 Flow-Attention on DAGs
For every DAG $\mathcal{D}\in\Delta$ with $\mathcal{D}\notin\mathcal{T}$ (i.e., a true DAG) and its computation tree $\gamma(\mathcal{D})$, there exists a flow-attentional directed acyclic GNN $\mathcal{A}_{\text{flow}}$ such that:
@@
for the graph $\mathcal{D}$ and its computation tree $\gamma(\mathcal{D})$, respectively. Since $\mathcal{D}$ is not a tree, there is at least one node $j$ with more than one successor. Hence, there is a predecessor $j$ for which $\beta_{ij}<1$. We then need to show that $\bm{h}^{\prime}_{i}\neq\tilde{\bm{h}}^{\prime}_{i}$:
@@ 5.3 FlowDAGNN
We propose a flow-attentional version of the attention-based DAGNN (Thost & Chen, 2021). A naive approach would be to simply replace the attention weights $\alpha_{ij}$ with flow attention weights $\beta_{ij}$. Due to the sequential nature of directed acyclic GNNs, the computation of the flow attention weight $\beta_{ij}$ in a DAG only depends on the node $i$ and all its ancestors. However, in contrast to standard attention weights, flow attention weights are forward-directed, which means that they should also be conditioned on all descendants of the node $i$. Analogously, the electrical current splitting up from one node into multiple branches depends on the whole branch and not only on the first node in each branch. In App. LABEL:app:example, we give a simple example of an electronic circuit illustrating this situation.
Hence, we construct a FlowDAGNN layer from two sublayers (see Fig. 2c). In the first sublayer (we call it the reverse pass), we invert all edges of the DAG $\mathcal{D}$ and apply a standard DAGNN layer to the reverse DAG $\tilde{\mathcal{D}}$. This is equivalent to performing the aggregation over all successor nodes in the original DAG $\mathcal{D}$ instead of over all predecessors: In flow attention, attention scores are normalized across outgoing neighbors, guiding a model to learn how to distribute a node’s message among them. This is especially useful on flow graphs governed by Kirchhoff’s first law, such as power grids, electronic circuits, or traffic networks. In contrast, other types of graph datasets, such as citation graphs or social networks, do not have an inherent structure that requires an information distribution across outgoing nodes. For example, there may be nodes that broadcast information to many adjacent nodes, e.g., influencers in social networks (Chen et al., 2009). Furthermore, as pointed out in the Introduction, some non-isomorphic graph structures might be equivalent in the context of informational graphs (e.g., computational graphs) and therefore should be mapped to the same representation, a consistency not guaranteed by a flow-attentional GNN due to Theorem 5.4. We therefore expect FlowGNNs to be less effective on these types of graph datasets. On the other hand, the flow-attention mechanism has the advantage over standard attention that it can distinguish between node multisets with the same distribution (Lemma 5.2). This could lead to performance improvements on graphs where certain node features occur repeatedly, which is often the case in flow graph applications (e.g., multiple identical components in electronic circuits), but can also be the case outside of the flow graph context (e.g., multiple identical atoms in molecules). In this study, we focus on evaluating the effectiveness of the flow attention mechanism on flow graph datasets, including power grids and electronic circuits. However, additional experimental results on standard graph datasets can be found in App. G.3.